# Hibou: A Family of Foundational Vision Transformers for Pathology

### v1

@@ Abstract
This paper introduces the Hibou family of foundational vision transformers for pathology, leveraging the DINOv2 framework to pretrain two model variants, Hibou-B and Hibou-L, on a proprietary dataset of over 1 million whole slide images (WSIs) representing diverse tissue types and staining techniques. Our pretrained models demonstrate superior performance on both patch-level and slide-level benchmarks, surpassing existing state-of-the-art methods. Notably, Hibou-L achieves the highest average accuracy across multiple benchmark datasets. To support further research and application in the field, we have open-sourced the Hibou-B model, which can be accessed at https://github.com/HistAI/hibou.
@@ 4.1 Patch-level benchmarks
Table 1: Linear probing benchmarks reporting top-1 accuracy. *Metrics for Virchow and RudolfV are derived from the respective papers, as these models are not open-sourced [7, 8].
Dataset Phikon [6] Kaiko-B8 [10] Virchow [7] RudolfV [8] Prov-GigaPath [12] Hibou-B Hibou-L
CRC-100K 0.917 0.949 0.968* 0.973* 0.968 0.955 0.966
PCAM 0.916 0.919 0.933* 0.944* 0.947 0.946 0.943
MHIST 0.791 0.832 0.834* 0.821* 0.839 0.812 0.849
MSI-CRC 0.750 0.786 - 0.755* 0.771 0.779 0.797
MSI-STAD 0.760 0.814 - 0.788* 0.784 0.797 0.825
TIL-DET 0.944 0.945 - 0.943* 0.939 0.942 0.943
AVG (1-3) 0.875 0.900 0.912 0.913 0.918 0.904 0.919
AVG (1-6) 0.846 0.874 - 0.871 0.875 0.872 0.887
@@ 5 Discussion and Future Work
Future work will focus on expanding our evaluation benchmarks to include additional subtyping tasks and new tasks like segmentation, which are critical for comprehensive histopathological analysis. Furthermore, we plan to investigate slide-level pretraining as this approach has the potential to improve the performance on WSI downstream tasks. Another promising direction for future research involves utilizing Hibou models as vision encoders in Large Vision-Language Models (LVLMs). These models integrate visual and textual data, enabling sophisticated interactions with histopathological slides. For instance, an LVLM could allow pathologists to query the model in natural language about specific features or abnormalities observed in a slide, receive detailed explanations, and even generate descriptive reports. This interactive capability could enhance diagnostic accuracy, streamline workflows, and facilitate a more intuitive and comprehensive analysis of histopathological data.

### v2

@@ Abstract
This paper introduces the Hibou family of foundational vision transformers for pathology, leveraging the DINOv2 framework to pretrain two model variants, Hibou-B and Hibou-L, on a proprietary dataset of over 1 million whole slide images (WSIs) representing diverse tissue types and staining techniques. Our pretrained models demonstrate superior performance on both patch-level and slide-level benchmarks, surpassing existing state-of-the-art methods. Notably, Hibou-L achieves the highest average accuracy across multiple benchmark datasets. To support further research and application in the field, we have open-sourced the Hibou models, which can be accessed at https://github.com/HistAI/hibou.
@@ 4.1 Patch-level benchmarks
Table 1: Linear probing benchmarks reporting top-1 accuracy. *Metrics for Virchow and RudolfV are derived from the respective papers. [7, 8].
Dataset Phikon [6] Kaiko-B8 [10] Virchow [7] RudolfV [8] Prov-GigaPath [12] H-optimus-0 [16] Hibou-B Hibou-L
CRC-100K 0.917 0.949 0.968* 0.973* 0.968 0.97 0.955 0.966
PCAM 0.916 0.919 0.933* 0.944* 0.947 0.942 0.946 0.953
MHIST 0.791 0.832 0.834* 0.821* 0.839 0.861 0.812 0.858
MSI-CRC 0.750 0.786 - 0.755* 0.771 0.767 0.779 0.793
MSI-STAD 0.760 0.814 - 0.788* 0.784 0.797 0.797 0.829
TIL-DET 0.944 0.945 - 0.943* 0.939 0.948 0.942 0.942
AVG (1-3) 0.875 0.900 0.912 0.913 0.918 0.924 0.904 0.926
AVG (1-6) 0.846 0.874 - 0.871 0.875 0.881 0.872 0.890
@@ 4.2 Slide-level benchmarks
4.3 Segmentation benchmarks
To evaluate the performance of the Hibou model on the segmentation task we employed a CellViT [18] framework and trained a segmentation model on a PanNuke dataset [19, 20]. PanNuke is an extensive, open-source pan-cancer histology dataset for nuclei instance segmentation and classification. We follow the training protocols described in [18], we train the model 3 times on different PanNuke splits and report averaged metrics. For more information on metrics and protocols check the original CellViT paper.
Table 3: Average PQ across the three PanNuke splits for each nuclear category. Metrics for CellViT256 and CellViT-SAM-H are from [18].
Model Neoplastic Epithelial Inflammatory Connective Dead
CellViT256 0.567 0.559 0.405 0.405 0.144
CellViT-SAM-H 0.581 0.583 0.417 0.423 0.149
CellViT-Hibou-L 0.582 0.591 0.426 0.425 0.185
Table 4: Average Precision, Recall, and F1 across the three PanNuke splits for each nuclear category. Metrics for CellViT256 and CellViT-SAM-H are from [18].
Model Neoplastic Epithelial Inflammatory Connective Dead
P R F1 P R F1 P R F1 P R F1 P R F1
CellViT256 0.69 0.70 0.69 0.68 0.71 0.70 0.59 0.58 0.58 0.53 0.51 0.52 0.39 0.35 0.37
CellViT-SAM-H 0.72 0.69 0.71 0.72 0.73 0.73 0.59 0.57 0.58 0.55 0.52 0.53 0.43 0.32 0.36
CellViT-Hibou-L 0.72 0.72 0.72 0.76 0.76 0.76 0.63 0.58 0.60 0.58 0.53 0.55 0.51 0.35 0.41
As demonstrated in Table 3 and Table 4 the CellViT-Hibou-L model, which uses the Hibou-L architecture as its backbone, consistently outperforms both the smaller ViT256 model, pretrained on pathological data [21], and a larger SAM-H [22] model, developed specifically for segmentation tasks but trained on natural images.
@@ 5 Discussion and Future Work
Future work will focus on expanding our evaluation benchmarks to include additional subtyping tasks, which are critical for comprehensive histopathological analysis. Furthermore, we plan to investigate slide-level pretraining as this approach has the potential to improve the performance on WSI downstream tasks. Another promising direction for future research involves utilizing Hibou models as vision encoders in Large Vision-Language Models (LVLMs). These models integrate visual and textual data, enabling sophisticated interactions with histopathological slides. For instance, an LVLM could allow pathologists to query the model in natural language about specific features or abnormalities observed in a slide, receive detailed explanations, and even generate descriptive reports. This interactive capability could enhance diagnostic accuracy, streamline workflows, and facilitate a more intuitive and comprehensive analysis of histopathological data.